Domino Company uses the aging of accounts receivable method to estimate uncollectible accounts expense. Domino began Year 2 with balances in Accounts Receivable and Allowance for Doubtful Accounts of $44,390 and $3,470, respectively. During the year, the company wrote off $2,640 in uncollectible accounts. In preparation for the company's Year 2 estimate, Domino prepared the following aging schedule: Number of days Receivables % Likely to be past due amount uncollectible Current $ 71,000 1% 0-30 27,000 5% 31-60 6,860 10% 61-90 3,520 25% Over 90 3,200 50% Total $ 111,580 What will Domino record as Uncollectible Accounts Expense for Year 2?
Answer #1

Un-adjusted Balance in Allowance account =Beginning balance -Write off

                   = 3470 - 2640

                  = $ 830

SCHEDULE OF ESTIMATED ALLOWANCE FOR DOUBTFUL ACCOUNT AT END
current 71000 1% 71000*1%= 710
0-30 27000 5% 1350
31-60 6860 10% 686
61-90 3520 25% 880
over 90 3200 50% 1600
ESTIMATED ALLOWANCE FOR DOUBTFUL ACCOUNT AT END 5226

Uncollectible Accounts Expense = Balance at end -unadjusted balance

             = 5226 - 830

             = $ 4396